OpenStudy (anonymous):

The number of chips of different colors in Vicky's bag is shown below: 5 blue chips 11 pink chips 9 white chips Vicky takes out a chip from the bag randomly without looking. She replaces the chip and then takes out another chip from the bag. What is the probability that Vicky takes out a blue chip in both draws? 5/25 * 5/25 = 25/625 5/25 + 5/25 = 10/25 5/25 * 4/24 = 20/600 5/25 + 4/24 = 220/600

OpenStudy (anonymous):

@Vocaloid

Vocaloid (vocaloid):

well, first let's find the probability of getting a blue chip number of blue chips/total number of chips = ?

OpenStudy (anonymous):

5/25 @Vocaloid

Vocaloid (vocaloid):

right, so to find the probability of 2 blue chips, we multiply (5/25) * (5/25) which is our answer

Vocaloid (vocaloid):

you can leave it as a fraction

OpenStudy (anonymous):

0.04 to 1/25

Vocaloid (vocaloid):

you don't have to reduce the fraction 5/25 * 5/25 = (5*5)/(25*25) = ?

OpenStudy (anonymous):

25/625 @Vocaloid